OpenStudy (anonymous):

Medal and fan If 25 mL of a .5 M NaOH solution were needed to reach the equivalence point of a 20 mL HCl solution, what is the original analyte concentration? Since I'm new to this side of chem, I don't know how to answer this question whatsoever.

OpenStudy (abmon98):

NaOH+HCl-->NaCl+H2O Number of moles=Concentration*volume N=0.5*25/1000=0.0125 moles Ratio of Moles of NaOH to HCl is 1:1 0.0125=Concentration*20/1000 20C/1000=0.0125 0.0125*1000/20=0.625 mol/dm^3

OpenStudy (anonymous):

Where do you get the 1000 from?

OpenStudy (abmon98):

Iam changing the volume unit from ml to dm^3

OpenStudy (anonymous):

Ohhh alright. Thankyou!

OpenStudy (abmon98):

your welcome :)